# Higher-Order Energy-Decreasing Exponential Time Differencing Runge–Kutta methods for Gradient Flows

### v1

@@ Abstract
We develop in this paper a general framework for constructing higher-order, unconditionally energy-stable exponential time differencing Runge–Kutta (ETDRK) methods applicable to a large class of gradient flows. Specifically, we identify a set of sufficient conditions that ETDRK schemes need to satisfy to maintain the original energy dissipation. Our analysis reveals that commonly used third-order and fourth-order ETDRK schemes fail to meet these conditions, and we construct new third-order ETDRK schemes which satisfy these conditions and thus guarantee the unconditional energy dissipation. We conduct extensive numerical experiments with these new schemes to verify their accuracy, stability, behavior under large time steps, long-term evolution, and adaptive time stepping strategy for various gradient flows.
@@ 1 Introduction
We consider a class of gradient flows written in the following general form
@@
where $\Omega$ is a bounded domain in $R^{d}$ ($d=1,2,3$), $T$ is a finite time, $G$ is a negative operator, $\mathcal{L}$ is a sectorial operator. The above system is energy dissipative with corresponding energy functional
@@
where $\mathcal{L}_{1/2}\mathcal{L}_{1/2}^{*}=\mathcal{L}_{1/2}^{*}\mathcal{L}_{1/2}=\mathcal{L}$ and $F^{\prime}=f$. This general form covers a wide range of gradient flows, such as the Allen–Cahn equation, the Cahn–Hilliard equation, the thin film model without slope selection also known as the molecular beam epitaxy (MBE) model, the phase-field crystal models and etc.
Since these gradient flows involve strong nonlinearities and often with high-order derivatives, it is difficult to design an efficient time discretization scheme which is able to accurately approximate their dynamics and steady states. In particular, it is challenging for numerical schemes to guarantee the energy decay which is intrinsic to all of these models. Ample numerical evidences indicate that non-physical oscillations may occur when the energy stability is violated. Furthermore, the establishment of unconditional energy stability enables the creation of more efficient algorithmic designs, including the development of adaptive time stepping strategies, where the selection of time steps is solely governed by considerations of accuracy.
There have been a large amount of studies concerning energy stable schemes for gradient flows, see, e.g. [9, 12, 10, 27, 30, 32, 28, 16, 18] and the references therein. In particular, the convex splitting technique [29, 25] is useful to preserve unconditional energy stability, but it leads to a nonlinear scheme and not easy to extend it to more general cases. The invariant energy quadratization (IEQ) method (see, e.g. [34, 35]) and the scalar auxiliary variable (SAV) method [4, 26] lead to linear schemes which can preserve different structures for a wide class of gradient flows. In particular, there have been a large number of works based on SAV methods, including SAV-BDF schemes [17, 3] and SAV-RK schemes [1, 31]. However, the energy stability achieved by the IEQ or SAV methods is based on a modified energy, not the original energy. We are not aware of any linear and higher than second-order scheme for (1.1) could be unconditional energy stable with respect to the original energy.
On the other hand, exponential time differencing Runge–Kutta (ETDRK) schemes offer many advantages over multistep methods In this paper, we consider exponential time differencing Runge–Kutta (ETDRK) methods for (1.1). The key idea of ETDRK schemes is to apply the Duhamel’s principle
@@
and approximate the implicit integral which contains the unknown $f(u(x,s))$ with an explicit RK method. The concept of exponential integrators has a long-standing history, tracing its origins back to the 1960s. There has been much literature related to using such methods to solve stiff problems, semilinear parabolic problems or gradient flows [6, 8, 11, 14, 5]. However, ETDRK schemes are usually not energy stable. It has been shown in [6, 7] that some first-order ETDRK schemes can be energy stable. More recently, it is shown in [11] that a second-order ETDRK scheme, with proper stabilization, is unconditional energy stable. However, no higher than second-order ETDRK scheme has proven to be energy stable.
This paper aims to introduce a comprehensive framework for developing unconditionally energy-stable ETDRK schemes tailored for various gradient flows. Our main contributions include
  • We have formulated a series of criteria for ensuring the unconditional energy stability of ETDRK schemes, irrespective of their order, see Theorem 2.1.
  • Our investigations reveal that the widely employed third-order and fourth-order ETDRK methods are not energy stable, and we have constructed new third-order ETDRK schemes that satisfy the criteria in Theorem 2.1, thereby achieving unconditional energy stability.
We are optimistic that the foundational framework proposed in this paper will pave the way for the creation of more advanced energy stable ETDRK schemes.
This paper is organized as follows. In Section 2, we present some background knowledge, then state and prove our main theorem. In Section 3, we apply the main theorem to various known ETDRK schemes and determine whether they are unconditional energy stable, in particular, we show that the commonly used third-order and fourth-order ETDRK schemes do not satisfy these conditions, and we construct new third-order ETDRK schemes which satisfy these conditions and thus are unconditional energy stable. In Section 4, we presents ample numerical experiments to validate our new third-order schemes, and to study how large time steps affect the solutions, how different time steps and stabilizers influence the energy evolution, we also present an example with adaptive time stepping. Some concluding remarks are given in the final section.
@@ 2 Energy decreasing ETDRK methods for gradient flows
In this section we first introduce some basic knowledge for gradient flows and some phase field models, convex splitting and the ETDRK methods. Then we prove our main theorem, i.e. arbitrary ETDRK schemes are unconditionally energy stable as long as the conditions are satisfied.
@@ 2.1 Phase-field models
[…]
@@
where $G$ is a negative operator, $\mathcal{L}$ is a symmetric sectorial operator, and $f(u)=F^{\prime}(u)$, where $F(u)$ is a nonlinear functional. Taking the inner product of (2.1) with $\mathcal{L}u+f(u)$, we find the energy dissipation law
@@
Some typical examples are:
  • the Allen–Cahn equation, $G=-1,\mathcal{L}=-\epsilon^{2}\Delta,f(u)=u^{3}-u$ and $F(u)=\frac{1}{4}(u^{2}-1)^{2}$;
@@
  • the MBE model without slope selection, $G=-1,\mathcal{L}=-\epsilon^{2}\Delta^{2}$ but $f(u)$ should be replaced by $f(\nabla u)=-\nabla\cdot(\frac{\nabla u}{1+|\nabla u|^{2}})$ and correspondingly $F(\nabla u)=-\frac{1}{2}\ln{(|\nabla u|^{2}+1)}$;
@@
We shall assume that $f$ is a Lipschitz continuous function with the Lipschitz constant $\beta$, i.e. we have
$\|f(u)-f(v)\|\leq\beta\|u-v\|,\;\forall\;u,\;v\in\mathcal{H}.$ (2.4)
This may appear to be somewhat restrictive since $f(u)$ in some popular phase-field models, such as Allen-Cahn, Cahn–Hilliard and phase field crystal, does not naturally satisfy the above Lipschitz assumption. However, as long as exact solutions of the problem in consideration are uniformly bounded, which is true for a large class of phase-field models, we can truncation $f(u)$ to quadratic growth as in [28, 11] so as to satisfy the Lipschitz assumption. In practice such modification does not affect convergence because of the boundedness of their solutions (cf. [21, 22], which offer solid analysis). Note that for the MBE model, the functions $F$ and $f$ are of $\nabla u$, which is different from other examples. The Lipschitz condition is satisfied in the sense
@@
where $L=\beta I+\mathcal{L}$ and $g=\beta I-f$ correspond to $E_{l}$ and $E_{n}$ respectively. In the computation, $Lu$ is treated implicitly and $g(u)$ is treated explicitly, which leads to linearly implicit schemes.
Here $\beta$ serves as a stabilization to enhance the dissipation of linear part, so as to bound the Lipschitz growing nonlinear term in the analysis. In fact, such modification or stabilization is necessary for the proof of energy decay, which is pointed out in the remark attached to the main theorem. However, the stabilization may cause significant time delay phenomena if a low-order scheme is used or a large time step is taken. We will study the effect in detail for higher-order schemes in Section 4. On the other hand, the stabilization is necessary to guarantee the MBP for Allen–Cahn equations [6]. Besides, [26] provides numerical evidences to illustrate that the stabilization can improve the numerical performance significantly.
@@ 2.2 ETDRK methods
where $\tau$ is the time step. For this scheme, the energy dissipation law and MBP for the Allen–Cahn equation particularly have been proved, see, e.g. [7]. The classical second order ETDRK (ETDRK2) for the gradient flow (1.1) reads
@@ 2.3 A general framework for energy stable ETDRK schemes
We present below a general framework for constructing energy stable ETDRK schemes. We start with a useful lemma whose proof is straightforward.
@@
Consider a positive-definite operator $L=\beta I+\mathcal{L}$, and let $f$ be an analytic function whose domain includes the spectrum of $L$, i.e. the values $\{f(\lambda_{i})\}_{i\in\mathcal{N}}$ exist, where $\{\lambda_{i}\}_{i\in\mathcal{N}}$ are the eigenvalues of $L$. Then, the eigenvalues of $f(L)$ are $\{f(\lambda_{i})\}_{i\in\mathcal{N}}$. Furthermore, if $f$ is a positive function, then $f(L)$ is also a positive-definite operator.
Hereafter, we say a matrix $\Delta$ is positive-definite if the eigenvalues of its symmetrizer $(\Delta+\Delta^{T})/2$ are all positive. The main result of this paper is stated below.
@@
where $G$ is a negative operator, $\mathcal{L}$ is a sectorial operator and $f$ is a Lipschitz continuous function with the Lipschitz constant $\beta$. The ETDRK schemes (2.12) unconditionally decreases the energy as long as the following determinant
@@
is positive-definite for all negative $z\in R$, where $I$ is the identity operator, $E_{L}=(1_{i\geq j})_{s\times s}$ is the lower triangular matrix with all nonzero entries equal to 1, and
@@
We first compute the difference of the energy and derive a key inequality.
@@
$\left(F(v)-F(u),1\right)$ $\leq(v-u,f(u))+\frac{\beta}{2}(v-u,v-u)$ (2.21)
$=-(v-u,g(u))+\beta(v-u,u)+\frac{\beta}{2}(v-u,v-u).$
@@
$E(v)-E(u)$ $\leq(v-u,Lv-g(u))-\beta(v-u,v-u)-\frac{1}{2}(v-u,\mathcal{L}(v-u))$ (2.23)
$\leq(v-u,Lv-g(u))-\frac{1}{2}(v-u,(\beta I+L)(v-u)).$
This vital inequality holds with the Lipschitz condition.
Now we focus on the ETDRK scheme (2.17), the first line of which is $v_{1}=u_{n}$, and the rest can be rewritten as the following system
@@
$\leq$ $\sum_{k=1}^{s}\left(v_{k+1}-v_{k},Lv_{k+1}-g(v_{k})\right)-\frac{1}{2}\sum_{k=1}^{s}(v_{k+1}-v_{k},(\beta I+L)(v_{k+1}-v_{k}))$
@@
$-\frac{\beta}{2}\sum_{k=1}^{s}(v_{k+1}-v_{k},v_{k+1}-v_{k})-\frac{1}{2\tau}\sum_{k=1}^{s}(v_{k+1}-v_{k},G^{-1}(\tau GL)(v_{k+1}-v_{k}))$
$=$ $\frac{1}{\tau}\sum_{k,j=1}^{s}(v_{k+1}-v_{k},G^{-1}\Delta_{kj}(\tau GL)(v_{j+1}-v_{j}))-\frac{\beta}{2}\sum_{k=1}^{s}(v_{k+1}-v_{k},v_{k+1}-v_{k}),$
where $\Delta(z)=zE_{L}+P^{-1}E_{L}-\frac{z}{2}I$ with $z=\tau GL$. Since the last term in the above is negative, and $G$ is a negative operator, we find that the discrete energy unconditionally decreases if $\Delta(z)$ is positive-definite for all negative $z\in R$.
@@
For the MBE model, recall that the nonlinear term is Lipschitz continuous as a function of $\nabla u$, and meanwhile the convex splitting of the energy is also different (Or see [33, 20] for more details), but all analysis in the proof can be carried out in the similar way. To keep the presentation short, we omit the proof.
@@
The inequality (2.23) in the proof plays a very important role. It is the only place where we apply the Lipschitz condition. Alternatively, we can also replace the Lipschitz condition by a convex splitting approach since
$E(v)-E(v)$ $\leq\left(v-u,\frac{\delta E_{l}}{\delta v}(v)-\frac{\delta E_{n}}{\delta u}(u)\right)$ (2.27)
@@
$=\left(v-u,Lv-g(u)\right)-\beta(v-u,v-u).$
The determinant of this version is slightly different but the rest of proof is the same.
@@
The positive-definiteness of $\frac{1}{2}(\Delta+\Delta^{T})$ is the same as the positive-definiteness of
@@
$\Delta^{\prime}$ is already symmetric and helps us in the understanding and subsequent proof regarding the positive-definiteness.

### v2

@@ Abstract
In this paper, we develop a general framework for constructing higher-order, unconditionally energy-stable exponential time differencing Runge–Kutta (ETDRK) methods applicable to a range of gradient flows. Specifically, we identify conditions sufficient for ETDRK schemes to maintain the original energy dissipation. Our analysis reveals that the widely employed third- and fourth-order ETDRK schemes fail to meet these conditions. To address this, we introduce new third-order ETDRK schemes, designed with appropriate stabilization, which satisfy these conditions and thus guarantee the unconditional energy decay property. We conduct extensive numerical experiments with these new schemes to verify their accuracy, stability, behavior under large time steps, long-term evolution, and adaptive time stepping strategy across various gradient flows. This study is the first to examine the unconditional energy stability of high-order ETDRK methods, and we are optimistic that our framework will enable the development of ETDRK schemes beyond the third order that are unconditionally energy stable.
@@ 1 Introduction
We consider a class of gradient flows written in the following general form:
@@
where $\Omega$ is a bounded domain in $R^{d}$ ($d=1,2,3$), $T$ is a finite time, $G$ is a nonpositive operator, and $\mathcal{L}$ is a positive definite operator. The above system is energy dissipative with the corresponding energy functional
@@
where $F$ is such that $F^{\prime}=f$. This general form covers a wide range of gradient flows, such as the Allen–Cahn equation, the Cahn–Hilliard equation, the thin film model without slope selection also known as the molecular beam epitaxy (MBE) model, the phase-field crystal models, etc.
Since these gradient flows involve strong nonlinearities and often high-order derivatives, it is difficult to design an efficient time discretization scheme which is able to accurately approximate their dynamics and steady states. In particular, it is challenging for numerical schemes to guarantee the energy decay which is intrinsic to all of these models. Ample numerical evidence indicates that non-physical oscillations may occur when the energy stability is violated. Furthermore, the establishment of unconditional energy stability enables the creation of more efficient algorithmic designs, including the development of adaptive time stepping strategies, where the selection of time steps is solely governed by considerations of accuracy.
There have been a large number of studies concerning energy-stable schemes for gradient flows, see, e.g. [9, 12, 10, 27, 30, 32, 28, 16, 18] and the references therein. In particular, the convex splitting technique [29, 25] is useful to preserve unconditional energy stability, but it leads to a nonlinear scheme and is not easy to extend to more general cases. The invariant energy quadratization (IEQ) method (see, e.g., [34, 35]) and the scalar auxiliary variable (SAV) method [4, 26] lead to linear schemes that can preserve different structures for a wide class of gradient flows. In particular, there have been a large number of works based on SAV methods, including SAV-BDF schemes [17, 3] and SAV-RK schemes [1, 31]. However, the energy stability achieved by the IEQ or SAV methods is based on a modified energy, not the original energy. We are not aware of any linear and higher than second-order for (1.1) which is unconditional energy stable with respect to the original energy.
On the other hand, exponential time differencing Runge–Kutta (ETDRK) schemes offer many advantages over multistep methods. In this paper, we consider exponential time differencing Runge–Kutta (ETDRK) methods for (1.1). The key idea of ETDRK schemes is to apply the Duhamel’s principle
@@
and approximate the implicit integral which contains the unknown $f(u(x,s))$ with an explicit RK method. The concept of exponential integrators has a long-standing history, tracing its origins back to the 1960s. There has been much literature related to using such methods to solve stiff problems, semilinear parabolic problems or gradient flows [6, 8, 11, 14, 5]. However, ETDRK schemes are usually not energy stable. It has been shown in [6, 7] that some first-order ETDRK schemes can be energy stable. More recently, it is shown in [11] that a second-order ETDRK scheme, with proper stabilization, is unconditionally energy stable. However, no ETDRK scheme higher than second-order has proven to be energy stable.
This paper aims to introduce a comprehensive framework for developing unconditionally energy-stable ETDRK schemes tailored for various gradient flows. We have formulated a series of criteria for ensuring the unconditional energy stability of ETDRK schemes of arbitrary order. Our investigations reveal that the widely employed third-order and fourth-order ETDRK methods fall short of meeting these criteria. In response, we have designed innovative third-order ETDRK schemes that adhere to these standards, thereby achieving unconditional energy stability. To our knowledge, this work pioneers the establishment of sufficient conditions for high-order ETDRK methods to attain unconditional energy stability. We are optimistic that this foundational framework will pave the way for the creation of advanced ETDRK schemes surpassing the third order in terms of unconditional energy stability.
This paper is organized as follows. In Section 2, we present some background knowledge, and then state and prove our main theorem. In Section 3, we apply the main theorem to various known ETDRK schemes and determine whether they are unconditionally energy-stable, in particular, we show that the commonly used third-order and fourth-order ETDRK schemes do not satisfy these conditions, and we construct new third-order ETDRK schemes that satisfy these conditions and thus are unconditional energy stable. In Section 4, we present ample numerical experiments to validate our new third-order schemes, and study how large time steps affect the solutions, how different time steps and stabilizers influence the energy evolution. We also present an example with adaptive time stepping to show the application of this scheme. Some concluding remarks are given in the final section.
@@ 2 Energy decreasing ETDRK methods for gradient flows
In this section we first introduce some basic knowledge of gradient flows, phase field models, convex splitting and the ETDRK methods. Then we prove our main theorem, i.e. arbitrary ETDRK schemes are unconditionally energy stable as long as the conditions are satisfied.
@@ 2.1 Phase-field models
We consider the general form of gradient flows in a Hilbert space $\mathcal{H}$:
@@
where $G$ is a nonpositive operator, $\mathcal{L}$ is a positive definite operator, and $f(u)=F^{\prime}(u)$, where $F(u)$ is a nonlinear functional. In the following text, we denote the inner product by (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). Taking the inner product of (2.1) with $\mathcal{L}u+f(u)$, we find the energy dissipation law:
@@
Some typical examples are provided below:
  • the Allen–Cahn equation, $G=-I,\mathcal{L}=-\epsilon^{2}\Delta,f(u)=u^{3}-u$ and $F(u)=\frac{1}{4}(u^{2}-1)^{2}$;
@@
  • the MBE model without slope selection, $G=-I,\mathcal{L}=-\epsilon^{2}\Delta^{2}$ but $f(u)$ should be replaced by $f(\nabla u)=-\nabla\cdot(\frac{\nabla u}{1+|\nabla u|^{2}})$, and correspondingly $F(\nabla u)=-\frac{1}{2}\ln{(|\nabla u|^{2}+1)}$;
@@
We shall assume that $f$ is a Lipschitz continuous function with the Lipschitz constant $C_{L}$, i.e. we have
$\|f(u)-f(v)\|\leq C_{L}\|u-v\|,\;\forall\;u,\;v\in\mathcal{H}.$ (2.4)
For the MBE model, the functions $F$ and $f$ depend only on $\nabla u$, which is different from other examples. The Lipschitz condition is satisfied in the sense
@@
Notice that for the Cahn–Hilliard equation and the phase field crystal model, $f(u)$ does not naturally satisfy the Lipschitz assumption on nonlinearity, but we can truncate $f(u)$ to quadratic growth as in [28, 11] so as to satisfy the Lipschitz assumption. In practice such a modification never affects anything because of the boundedness of their solutions, see [21, 22], which offer solid analysis without Lipschitz assumptions.
@@
where $L=\beta I+\mathcal{L}$ and $g=\beta I-f$ correspond to $E_{l}$ and $E_{n}$, respectively. In the computation, $Lu$ is treated implicitly and $g(u)$ is treated explicitly, which leads to linearly implicit schemes.
Here $\beta$ serves as a stabilization to enhance the dissipation of linear part, so as to bound the Lipschitz growing nonlinear term in the analysis. In fact, such modification or stabilization is necessary for the proof of energy decay, which is pointed out in the remark attached to the main theorem. However, the stabilization may cause significant time delay phenomena if a low-order scheme is used or a large time step is taken. We will study the effect in detail for higher-order schemes in Section 4. On the other hand, the stabilization is necessary to guarantee the maximum bound property (MBP) for Allen–Cahn equations [6]. Besides, [26] provides numerical evidence to illustrate that the stabilization can significantly improve the numerical performance.
@@ 2.2 ETDRK methods
where $\tau$ is the time step. For this scheme, the energy dissipation law and MBP for the Allen–Cahn equation have been proved, see, e.g., [7]. The classical second-order ETDRK (ETDRK2) for the gradient flow (1.1) reads
@@ 2.3 A general framework for energy stable ETDRK schemes
We present below a general framework for constructing energy-stable ETDRK schemes. We start with a useful lemma whose proof is straightforward.
@@
Consider a positive-definite operator $L=\beta I+\mathcal{L}$, and let $f$ be an analytic function whose domain includes the spectrum of $L$, i.e., the values $\{f(\lambda_{i})\}_{i\in\mathcal{N}}$ exist, where $\{\lambda_{i}\}_{i\in\mathcal{N}}$ are the eigenvalues of $L$. Then, the eigenvalues of $f(L)$ are $\{f(\lambda_{i})\}_{i\in\mathcal{N}}$. Furthermore, if $f$ is a positive function, then $f(L)$ is also a positive-definite operator.
Hereafter, we say a matrix $\Delta$ is positive-definite if the eigenvalues of its symmetrizer $(\Delta+\Delta^{T})/2$ are all positive.
@@
where $G$ is a negative operator, $\mathcal{L}$ is a sectorial operator and $f$ is a Lipschitz continuous function with the Lipschitz constant $C_{L}$. The ETDRK schemes (2.12) with the stabilizer $\beta\geq C_{L}$ unconditionally decreases the energy as long as the following determinant
@@
is positive-definite for all negative $z\in\mathbb{R}$, where $I$ is the identity operator, $E_{L}=(1_{i\geq j})_{s\times s}$ is the lower triangular matrix with all nonzero entries equal to 1, and
@@
We first compute the difference in the energy and derive a key inequality.
@@
$\left(F(v)-F(u),1\right)$ $\leq(v-u,f(u))+\frac{C_{L}}{2}(v-u,v-u)$ (2.21)
$=-(v-u,g(u))+\beta(v-u,u)+\frac{C_{L}}{2}(v-u,v-u).$
@@
$E(v)-E(u)\leq(v-u,Lv-g(u))-\frac{1}{2}(v-u,L(v-u))-\frac{\beta-C_{L}}{2}(v-u,v-u).$ (2.23)
This vital inequality holds with the Lipschitz condition and if we take $\beta\geq C_{L}$, the second term is naturally non-positive.
Now we focus on the ETDRK scheme (2.17), the first line of which is $v_{1}=u_{n}$, and the rest can be rewritten as the following system:
@@
$\leq$ $\sum_{k=1}^{s}\left(v_{k+1}-v_{k},Lv_{k+1}-g(v_{k})\right)-\frac{1}{2}\sum_{k=1}^{s}(v_{k+1}-v_{k},L(v_{k+1}-v_{k}))-\frac{\beta-C_{L}}{2}\sum_{k=1}^{s}\|v_{k+1}-v_{k}\|^{2}$
@@
$-\frac{1}{2\tau}\sum_{k=1}^{s}(v_{k+1}-v_{k},G^{-1}(\tau GL)(v_{k+1}-v_{k}))-\frac{\beta-C_{L}}{2}\sum_{k=1}^{s}\|v_{k+1}-v_{k}\|^{2}$
$=$ $\frac{1}{\tau}\sum_{k,j=1}^{s}(v_{k+1}-v_{k},G^{-1}\Delta_{kj}(\tau GL)(v_{j+1}-v_{j}))-\frac{\beta-C_{L}}{2}\sum_{k=1}^{s}\|v_{k+1}-v_{k}\|^{2},$
where $\Delta(z)=zE_{L}+P^{-1}E_{L}-\frac{z}{2}I$ with $z=\tau GL$. Since $G$ is a negative operator and $\beta\geq C_{L}$, the discrete energy unconditionally decreases if $\Delta(z)$ is positive-definite for all negative $z\in R$.
@@
For the MBE model, recall that the nonlinear term is Lipschitz continuous as a function of $\nabla u$, and meanwhile the convex splitting of the energy is also different (see [33, 20] for more details). However, all analysis in the proof can be carried out in the similar way. To keep the presentation short, we omit the proof.
@@
The inequality (2.23) in the proof plays a very important role. It is the only place where we apply the Lipschitz condition. Alternatively, we can also replace the Lipschitz condition with a convex splitting approach since
$E(v)-E(u)$ $\leq\left(v-u,\frac{\delta E_{l}}{\delta v}(v)-\frac{\delta E_{n}}{\delta u}(u)\right)$ (2.27)
@@
$=\left(v-u,Lv-g(u)\right).$
The determinant of this version is slightly different, while the rest of the proof is the same.
@@
The positive-definiteness of $\frac{1}{2}(\Delta+\Delta^{T})$ is equivalent to the positive-definiteness of
@@
$\Delta^{\prime}$ is already symmetric and helps us in the understanding and subsequent proof regarding positive-definiteness.